# NutePrune: Efficient Progressive Pruning with Numerous Teachers for Large Language Models

### v1

@@ 4.3 Ablation Study
Sparsity gap between teacher and student
During stage 1, the sparsity gap between teacher and student model is an important hyperparameter. We conduct experiments with various sparsity gaps and the results are presented in Table 8. A 10% gap is deemed appropriate to prevent a gap that is too small, as it may result in insufficient guidance, or a gap that is too large, which would also toughing distillation.
Sparsity Gap Avg.(%)
5% 53.62
10% (ours) 53.94
20% 53.04
@@
Snapshot interval of teachers
When taking snapshots of students as teachers, it is preferable to save as many teachers as possible to facilitate more comprehensive training. However, the additional CPU storage for these teachers incurs extra costs. As demonstrated in Table 9, selecting an interval of 1% leads to significant improvement over the 10% interval, and the associated extra storage is acceptable, which means we use 10 teachers for 20% sparsity and 40 teachers for 50% sparsity.
@@ 5 Conclusion
In this work, we propose NutePrune as a novel efficient pregressive structured pruning method for LLMs. By minimizing the cost of KD, NutePrune aligns closely with the memory and time requirements of conventional training, allowing the utilization of numerous teachers for distillation. This approach mitigates the capacity gap between teacher and student, preserving the quality of the distillation process. We showcase the effectiveness of NutePrune, particularly at high sparsity levels, across various tasks and metrics for LLMs. The findings of this work contribute to the advancement of structured pruning techniques for LLMs, particularly in resource-constrained scenarios.
@@ 6 Limitations
We acknowledge the main limitation of this work is that we only evaluate our methods on LLaMA with 7B and 13B due to limited computation resources. And other model families, such as LLaMA-v2, are not included due to the absence of results for comparison from state-of-the-art methods. Recent work Ma et al. (2023); Xia et al. (2023) proves that using extensive data for post-training could substantially enhance the performance, but it comes with a substantial increase in computational costs. We target on pruning on resource-constraint scenarios and leave pruning with extensive data for future work.
Appendix A Zero-shot Performance with Newer Version
@@
Ratio Tune Method BoolQ PIQA HellaSwag WinoGrande ARC-e ARC-c OBQA Avg.
@@
Appendix B Pruned Structure
To gain insights into the pruned model, we present a detailed overview of the pruned structure at sparsity levels of 20% and 50%. The original hidden dimension is 4096, with a number of heads set at 32 and an intermediate dimension of 11008. Tables 11 and 12 reveal several observations. Notably, NutePrune tends to avoid pruning the hidden dimension, which aligns with the observation that pruning it may result in significant performance degradation Ma et al. (2023). Regarding heads and intermediate dimensions, NutePrune tends to prune the the last few layers. This observation differs from LLM-Pruner, which asserts the importance of the last layers. Further analysis of this phenomenon is left for future work.
@@ Appendix B Pruned Structure
Appendix C Generated Examples
@@
Model
Generated Sentences
w/o Tune
Neural Sequential Model, especially transformers, are used to model the human brain. The model is based on the assumption that the brain is a system of interconnected modules, each of which is responsible for a specific function. The modules are connected by synapses, which are the synaptic connections between neurons. The synapsis are responsible for the transfer of information between the modules. The information is transferred in the form of a signal, which is a combination of the input signal and the output signal. The input signal is the signal that is received by the module, while the output is the output of the module. The output signal is then passed on to the next …
@@
w/ Tune
Neural Sequential Model, especially transformers, are used to model the human brain. The model is a neural network that is trained to recognize patterns in data. The network is trained by feeding it data and then testing it on new data. This is a form of machine learning. The neural network is a type of artificial neural network. The term neural network refers to a group of neurons that are connected together. The neurons are connected by synapses. The synapsis are the connections between the neurons. The connections are called weights. The weights are used by the network to determine how much influence each neuron has on the other neurons in the network …
@@
w/ Tune + Post FT
Neural Sequential Model, especially transformers, is a popular model for deep learning. It is a model that can be trained on large datasets. It is used for tasks such as image classification, natural language processing, and other tasks. It is an effective model for tasks such as image classification and natural language processing. It is also used for tasks like natural language understanding and natural film fiction. It is able to learn from large datasets and can be trained for various tasks. It can be used for tasks that require a large amount of data. It is capable of learning from large datasets. It is used to train neural networks for tasks such image classification …

### v2

@@ 4.3 Ablation Study
Sparsity gap and interval of teachers
During stage 1, the sparsity gap between teacher and student model is an important hyperparameter. As shown in Table 12, a 10% gap is deemed appropriate to prevent a gap that is too small, as it may result in insufficient guidance, or a gap that is too large, which would toughing distillation. And when taking snapshots of students as teachers, it is preferable to save as many teachers as possible to facilitate more comprehensive training. However, it comes with extra costs. As demonstrated in Table 13, selecting an interval of 1% leads to significant improvement over the 10% interval, and the associated extra storage is acceptable.
Sparsity Gap 5% 10% 20%
Avg.(%) 53.62 53.94 53.04
@@ 5 Conclusion
In this work, we propose NutePrune as a novel efficient progressive structured pruning method for LLMs. Our well-designed techniques minimize the memory cost of KD, enabling NutePrune to utilize numerous teachers to mitigate the capacity gap between teacher and student and improve the quality of distillation. We show the effectiveness of NutePrune across various base models on diverse metrics. This work contributes to structured pruning techniques for LLMs, particularly in resource-constrained scenarios.
@@ 6 Limitations
Recent work Ma et al. (2023); Xia et al. (2023) proves that using extensive data for post-training could substantially enhance the performance, but it comes with a substantial increase in computational costs. We target on pruning on resource-constraint scenarios and leave pruning with extensive data for future work.
Appendix A Detailed $L_{0}$ Regularization
To get the learnable masks $\mathbf{z}$, $L_{0}$ regularization introduces a sampling strategy as augmentation during training. $\mathbf{z}$ is a real number in $[0,1]$ and is obtained from:
$\mathbf{u}$ $\sim U(0,1)$ (10)
$\tilde{\mathbf{s}}$ $=\mathbf{s}\times(r-l)+l$
$\mathbf{z}$ $=\min(1,\max(0,\tilde{\mathbf{s}})),$
Appendix B Zero-shot Performance with Newer Version
@@
[…]
@@
Appendix C Pruning at Higher Sparsity
To demonstrate the effectiveness of NutePrune at higher sparsity, we conducted experiments at 70% sparsity in Table 14.
[…]
0% LLaMA-7B 5.68 75.11 79.16 76.21 69.85 75.29 44.71 44.40 66.39
70% LLM-Pruner 56.33 47.28 60.83 31.66 50.75 39.56 24.83 28.80 40.53
NutePrune 34.30 62.08 62.30 39.43 51.46 42.17 26.19 30.20 44.83
Table 15: Performance (%) of the compressed LLaMA-7B models at 70% sparsity.
Appendix D Pruned Structure
To gain insights into the pruned model, we present a detailed overview of the pruned structure at sparsity levels of 20% and 50%. The original hidden dimension is 4096, with a number of heads set at 32 and an intermediate dimension of 11008. Tables 16 and 17 reveal several observations. Notably, NutePrune tends to avoid pruning the hidden dimension, which aligns with the observation that pruning it may result in significant performance degradation Ma et al. (2023). Regarding heads and intermediate dimensions, NutePrune tends to prune the the last few layers. This observation differs from LLM-Pruner, which asserts the importance of the last layers. Further analysis of this phenomenon is left for future work.
@@ Appendix D Pruned Structure
Appendix E Generated Examples
@@
Model Generated Sentences
w/o Tune Neural Sequential Model, especially transformers, are used to model the human brain. The model is based on the assumption that the brain is a system of interconnected modules, each of which is responsible for a specific function. The modules are connected by synapses, which are the synaptic connections between neurons. The synapsis are responsible for the transfer of information between the modules. The information is transferred in the form of a signal, which is a combination of the input signal and the output signal. The input signal is the signal that is received by the module, while the output is the output of the module. The output signal is then passed on to the next …
@@
w/ Tune Neural Sequential Model, especially transformers, are used to model the human brain. The model is a neural network that is trained to recognize patterns in data. The network is trained by feeding it data and then testing it on new data. This is a form of machine learning. The neural network is a type of artificial neural network. The term neural network refers to a group of neurons that are connected together. The neurons are connected by synapses. The synapsis are the connections between the neurons. The connections are called weights. The weights are used by the network to determine how much influence each neuron has on the other neurons in the network …
@@
w/ Tune + Post FT Neural Sequential Model, especially transformers, is a popular model for deep learning. It is a model that can be trained on large datasets. It is used for tasks such as image classification, natural language processing, and other tasks. It is an effective model for tasks such as image classification and natural language processing. It is also used for tasks like natural language understanding and natural film fiction. It is able to learn from large datasets and can be trained for various tasks. It can be used for tasks that require a large amount of data. It is capable of learning from large datasets. It is used to train neural networks for tasks such image classification …